Write a functional programme in Python for the following task:   Consider the quantum mechanical problem of a particle encountering a potential step of height V. The particle with wave number enters from the left and meets the potential step at 0. If the kinetic energy Eof the particle is larger than V, it can either pass the step and continue with a smaller wave number or be reflected keeping its kinetic energy. The formulae for the probability of transmission (T) or reflection (R) are given as (ki + k2)2 k1 + k2 Write a function trprob) that expects as single input number the kinetic energy, E, in units of eV and calculates the transmission and reflection probabilities using . an electron with mass m-511 × 103 eV C-2, meeting a potential step of height V-9 eV . Take as data for the formula 4.135667 eV s Safeguard (see accompanying PDF file for details) your function against (for this problem) not permitted kinetic energy values E < V Return the transmission and reflection probabilities as a tuple (in that order)
Answer #1

i have written program in python with output as well

Editor-Canopy File Edit View Search Run Tools Window Help free pertide.py tprob.py 1 import math 2 def tprob(E): This function returns the transmission and reflection probability for given energy value E value of the entered number V = 9.0 m5.11E5 hbar4.135667 if EV: k1math.sqrt (2.8*m*E)/hbar k2math.sqrt (2.m(E-V))/hbar dermath.pow( (k1+k2),2) neri = 4.0*k1rk2 ner2 math.pow( (k1-k2),2) trner1/der rr = ner2/der return(transmission coefficient, tr, reflection coefficient 10 14 15 16 rr) else: 18 19 20 21 # output: for E = 6 eV 22 print (tprob(16.0)) 23 24#output : for E- 16 eV 25 print (tprob(6.0)) 26 return (no value as E<v) Cursor pos 5:17 Pvthon 2 D:/Readings and books/project/tprob.py 8-53 PM O Type here to searchEditor - Canopy File Edit View Search Run Tools Window Help free partide.py trob.ру 1 import math 2 def tprob(E): This function returns the transmission and reflection probability for given energy value E value of the entered number V 9.0 m5.11E5 În no [3]: %run value D: /Readings D: /Readings D: /Readings and books/project/tprob.py books/project/tprob.py books/project/tprob.py In no [4]: %run value and In no [5]: %run value and In [6]: %run D: /Readings and books/project/tprob.py transmission coefficient ,0.9584750922440659) In [7J: %run D: /Readings and books/project/tprob.py transmission coefficient , 0.9584750922440659, reflection coefficiet 0.041524907755934086) In [8]: %run D: /Readings and books/project/tprob.py transmission coefficient , 0.9584750922440659, reflection coefficiet no value as EV 0.041524907755934086) In [91 Cursor pos 3: 102 Pvthon 2 D:/Readings and books/project/tprob.py O Type here to search 8-52 PM 1/21/2019 6please rate it up thanks :)